### v1

# USTM: Unified Spatial and Temporal Modeling for Continuous Sign Language Recognition

Abstract
Continuous sign language recognition (CSLR) requires precise spatio-temporal modeling to accurately recognize sequences of gestures in videos. Existing frameworks often rely on CNN-based spatial backbones combined with temporal convolution or recurrent modules. These techniques fail in capturing fine-grained hand and facial cues and modeling long-range temporal dependencies. To address these limitations, we propose the Unified Spatio-Temporal Modeling (USTM) framework, a spatio-temporal encoder that effectively models complex patterns using a combination of a Swin Transformer backbone enhanced with lightweight temporal adapter with positional embeddings (TAPE). Our framework captures fine-grained spatial features alongside short and long-term temporal context, enabling robust sign language recognition from RGB videos without relying on multi-stream inputs or auxiliary modalities. Extensive experiments on benchmarked datasets including PHOENIX14, PHOENIX14T, and CSL-Daily demonstrate that USTM achieves state-of-the-art performance against RGB-based as well as multi-modal CSLR approaches, while maintaining competitive performance against multi-stream approaches. These results highlight the strength and efficacy of the USTM framework for CSLR. The code is available at https://github.com/gufranSabri/USTM
I Introduction
Sign language is a fully-fledged visual language used by deaf communities for communication. This language is characterized by a complex combination of hand motion, facial expression, and body posture [wadhawan2021sign]. Unlike spoken languages, which rely on sequential acoustic signals, sign languages are inherently multimodal and spatio-temporal, which encode information simultaneously through both manual and non-manual cues [sidig2021karsl]. Sign Language Recognition (SLR) aims to automatically recognize sign language gestures in videos and translate them into written or spoken language. SLR is important serving as a crucial step toward integrating deaf people in the community. It can be broadly categorized based on the recognized sign unit into isolated SLR (ISLR) and continuous SLR (CSLR). ISLR systems focus on recognizing individual signs that are already segmented, whereas CSLR systems aim to recognize sign language sentences in which signs appear fluidly in sequence, much like natural speech. CSLR is more challenging than ISLR due to the lack of temporal information that marks the precise start and end of each sign in the sentence making it a weakly supervised learning problem [luqman2023arabsign]. This ambiguity in temporal boundaries, coupled with high inter-signer and intra-signer variability, makes CSLR more challenging than its isolated counterpart.
Several approaches have been proposed in the literature for CSLR. These approaches can be categorized into single-modality and multi-modality methods. The majority of recent CSLR studies [cheng2020fully, min2021visual, hao2021self, xie2023multi, hu2023self, cui2023spatial, alyami2025swin] predominantly utilize RGB-based modality. Cheng et al. [cheng2020fully] proposed a fully convolutional network with gloss feature enhancement to improve alignment between visual features and gloss sequences. To better capture local motion cues, Xie et al. [xie2023multi] introduced a multi-scale temporal similarity fusion network with position-aware convolutions. VAC loss [min2021visual] further improved alignment by enforcing visual consistency between encoder and decoder features. Attention and distillation-based methods, such as SEN [hu2023self] and SMKD [hao2021self], emphasized key regions (e.g., hands and face) and reduced prediction spikes. Temporal Lift Pooling [hu2022temporal] and CorrNet [hu2023continuous] techniques tackled temporal downsampling and cross-frame correlation modeling. To address scalability on longer videos, transformer-based models, such as ST-Transformer [cui2023spatial] and Swin-MSTP [alyami2025swin], leveraged frame chunking and multiscale temporal modules, showing strong performance with RGB-only input.
Multi-modal and multi-stream approaches have also shown a strong promise in advancing CSLR by incorporating complementary information beyond raw RGB frames. Multi-modal models leverage additional cues, such as pose, hand crops, and optical flow, to enrich spatial and temporal representations. SignBERT [zhou2021signbert] incorporates hand-focused visual cues into a transformer backbone to enhance signer generalization, while C2SLR [zuo2022c2slr] employs keypoint-guided attention alongside semantic constraints. STMC [zhou2021spatial] and Deep Neural Fusion [cui2019deep] combine pose, RGB, and optical flow modalities to enhance temporal modeling. Modality alignment [li2022multi] and dual-stream networks [chen2022two] further improve CSLR performance by capturing complementary patterns across input types. Multi-stream systems, on the other hand, process the same video through multiple parallel pathways to capture diverse temporal features. AdaBrowse [hu2023adabrowse] introduces an adaptive selection mechanism to sample informative subsequences and resolutions via a lightweight policy network. Similarly, SlowFastSign [ahn2024slowfast] extends the SlowFast paradigm developed for action recognition with bidirectional fusion and spatial-temporal enhancement modules, achieving high performance without additional inference overhead. While these methods perform well on CSLR, they typically introduce higher computational cost and require modality-specific supervision, posing limitations for practical deployment.
Most RGB-based CSLR models adopt a pipeline consisting of a spatial encoder followed by a temporal encoder and an alignment module [alyami2024reviewing]. This sequential separation postpones temporal modeling to later stages which prevents early layers from jointly modeling the rich spatio-temporal dynamics crucial for recognizing subtle and context-dependent sign language cues. To address this issue, we propose a Unified Spatial and Temporal Model (USTM) for CSLR. The proposed framework tightly integrates spatial and temporal modeling early in the recognition pipeline. USTM uses a Swin transformer with a custom Temporal Adapter with Positional Embeddings (TAPE) integrated after each feature extraction stage of the proposed model. The TAPE adapter enables joint spatio-temporal feature extraction. This enhanced Swin transformer is followed by a temporal enhancement module that combines a Multi-Scale Temporal Convolution Network (MS-TCN) for capturing fine-grained local dynamics and a bi-directional LSTM to model long-range temporal dependencies. Then, a CTC-based alignment module is used to produce gloss-level outputs without requiring precise temporal annotations. Our contributions are four-fold: (1) evaluating the effectiveness of transformer-based backbones by comparing ViT and Swin for CSLR; (2) proposing TAPE, a temporal adapter designed to adapt image-based vision backbones for video understanding tasks; (3) conducting an extensive ablation study to investigate the individual contribution of each component within our framework and (4) achieving a new state-of-the-art (SOTA) performance on the PHOENIX14 [phoenix14] and PHOENIX14T [phoenix14t] datasets, while attaining competitive results on the CSL-Daily [csldaily] dataset under single-stream RGB-only conditions as well as in multi-modality and multi-stream settings.
The remaining paper is organized as follows: Section II provides a review of related works; Section III describes our proposed framework; Section IV details our experimental setup, results, and ablation studies; and Section V concludes the paper.
II Related Work
CSLR Approaches. Single-stream CSLR approaches process only RGB videos through a single pathway, aiming to capture spatial-temporal dependencies without leveraging additional modalities or parallel streams. Cheng et al. [cheng2020fully] proposed a fully convolutional framework with gloss feature enhancement to improve alignment. To better capture local motion, Xie et al. [xie2023multi] introduced a multi-scale temporal similarity fusion network using position-aware convolutions. VAC loss [min2021visual] enforces alignment consistency between modules, improving feature quality. Hu et al. [hu2023self] emphasized hand/face regions using spatial-temporal attention, while Hao et al. [hao2021self] used mutual distillation to reduce CTC spikes. Temporal Lift Pooling (TLP) [hu2022temporal] improves temporal downsampling, and CorrNet [hu2023continuous] models cross-frame motion via dynamic correlations. To handle long sequences, Cui et al. [cui2023spatial] proposed a spatial-temporal transformer using frame chunking. Wang et al. [wang2025continuous] proposed STNet, which strengthens frame-wise and multi-scale spatio-temporal representations through a spatial resonance module that captures global common spatial features across adjacent frames, together with a multi-temporal perception module. Similarly, Gan et al. [gan2024signgraph] proposed SignGraph, where sign sequences are represented as graphs to capture intra- and inter-frame cross-region interactions across multiple scales. Most recently, Alyami et al. [alyami2025swin] enhanced Swin Transformers with multiscale temporal modules for fine-grained motion modeling.
Multi-modal CSLR models leverage additional cues such as pose, optical flow, and hand/face features to enhance recognition performance. Cui et al. [cui2019deep] combined CNN-RNN architectures with iterative RGB-flow fusion to better capture temporal dependencies. SignBERT [zhou2021signbert] integrates hand crops with transformer-based models to improve generalization across different signers. Zhou et al. [zhou2021spatial] proposed a spatial-temporal multi-cue network that jointly models multiple visual cues through dedicated spatial and temporal modules. C2SLR [zuo2022c2slr] employs keypoint-guided spatial attention and sentence embedding consistency constraints to enhance both visual and sequential modules. TwoStream-SLR [chen2022two] and CVT-SLR [zheng2023cvt] explore dual-stream and contrastive visual-textual alignment strategies, combining raw RGB, keypoint sequences, and pretrained language modules for more robust cross-modal representations. Multi-channel transformers [camgoz2020multi] explicitly model intra- and inter-channel relationships of different articulators, while temporal deformable convolutional sequence learning [papadimitriou2020multimodal] fuses multiple spatio-temporal streams for end-to-end recognition. Finally, frameworks such as SAM-SLR [jiang2021skeleton] incorporate graph-based skeleton modeling alongside RGB and depth cues.
Figure 1: The USTM framework. Spatio-temporal features are extracted using a Swin Transformer backbone with TAPE adapters and subsequently processed by MSTCN and BiLSTM modules in the temporal enhancement stage.
Multi-stream models are designed with multiple parallel pathways to process the same type of input modality, enabling the extraction of complementary features. In the context of CSLR, these models have demonstrated considerable improvements in capturing both spatial and temporal information. For instance, the SlowFast network [ahn2024slowfast] employs a dual-pathway architecture, where one pathway operates at a high temporal resolution to capture dynamic motion, while the other focuses on spatial details, such as hand shapes and facial expressions. AdaBrowse [hu2023adabrowse] introduces a lightweight adaptive mechanism that selectively processes the most informative temporal subsequences and dynamically adjusts input resolution. Similarly, cross-modal augmentation strategies [pu2020boosting] generate pseudo video-text pairs to enhance alignment between video features and textual labels. Beyond RGB inputs, keypoint-based multi-stream models, such as MSKA-SLR [guan2405multi], leverage skeletal information to reduce background noise sensitivity and computational complexity. More recently, a graph-based multi-stream approach, namely MSD-GCN [rastgoo2025multi], integrated adaptive motion-aware attention, signal-enhanced feature extraction, and self-supervised pretraining to capture fine-grained motion patterns and long-term dependencies.
Figure 2: Architecture of TAPE adapter.
Transformer Architectures for Video Understanding. Vision Transformers (ViTs) [dosovitskiy2020image] introduced a new paradigm in visual learning by applying self-attention over image patches to effectively capture long-range dependencies without convolution. While ViT showed strong performance on large-scale datasets, follow-up models, such as DeiT [touvron2021training], T2T-ViT [yuan2021tokens], and PVT [wang2021pyramid], improved performance in data-constrained settings via distillation, better tokenization, and hierarchical design. Later variants like Swin Transformer [liu2021swin] and CvT [wu2021cvt] introduced local inductive biases and convolutional priors, enabling ViTs to close the performance gap with CNNs on dense prediction tasks. These developments paved the way for transformer-based architectures in video understanding, where modeling both spatial and temporal dynamics is essential.
To extend ViTs to video, several models have been proposed to explicitly capture temporal dependencies. TimeSformer [bertasius2021space] factorized attention into spatial and temporal components, demonstrating competitive performance on action recognition tasks. ViViT [arnab2021vivit] examined various attention schemes, including joint and factorized attention with tubelet embeddings. MViT [fan2021multiscale] introduced multi-scale attention and pooling mechanisms, while Uniformer [li2023uniformer] unified convolution and self-attention for efficient spatial-temporal modeling. More recently, STAdapter [pan2022st] proposed injecting temporal reasoning into spatial backbones using lightweight temporal adapters. This approach offers a modular and parameter-efficient solution for adapting pretrained vision transformers to video-based tasks.
III Methodology
Given an input video $\mathcal{X}=\{x_{1},\dots,x_{n}\}$ consisting of $n$ frames, the objective of CSLR is to predict a sequence of $m$ glosses $\mathcal{Y}=\{y_{1},\dots,y_{m}\}$ representing the signs that appear in the video. Most existing CSLR frameworks follow a three-stage pipeline consisting of a spatial module $\mathcal{F}_{s}(\cdot)$ to extract visual features from each frame, a temporal module $\mathcal{F}_{t}(\cdot)$ to model motion and dependencies across frames, and an alignment module $\mathcal{F}_{a}(\cdot)$, typically trained with CTC loss, to generate the final gloss sequence. The pipeline can be formulated as follows:
$\mathcal{Y}=\mathcal{F}_{a}(\mathcal{F}_{t}(\mathcal{F}_{s}(\mathcal{X})))$
In contrast to this traditional way of processing spatio-temporal data, we adopt an integrated framework that departs from this strictly sequential design. The proposed framework USTM, consists of a stack of Spatio-Temporal Modules (STeMs) followed by a temporal enhancement stage that refines motion representations using MS-TCN and BiLSTM modules. Finally, a CTC-based alignment module decodes the output into gloss sequences. Figure 1 shows the USTM pipeline.
III-A Modeling Spatio-Temporal Features
To effectively model both spatial and temporal patterns in sign language videos, we propose a vision encoder consisting of a stack of STeM blocks. As illustrated in Figure 1, each STeM block consists of a Swin transformer block for spatial feature extraction, followed by a TAPE adapter for temporal modeling. The motivation for TAPE comes from the observation that temporal modeling in many CSLR architectures typically occurs after spatial features are pooled or flattened into 1D vectors, which limits the ability to capture fine-grained spatio-temporal interactions. TAPE addresses this limitation by applying temporal processing within the spatial feature extraction stage.
The TAPE adapter is a lightweight and pluggable module designed to enhance spatial features with short-term temporal context. As illustrated in Figure 2, the TAPE adapter begins with layer normalization followed by a down-projection to a lower-dimensional latent space. Learnable temporal positional embeddings are then added to encode frame order. Temporal modeling within TAPE is carried out through two components: Channel Mix (CMix) and Local Spatio-Temporal (LST) modules. The CMix component applies a $1\times 1\times 1$ convolution followed by batch normalization, allowing efficient recombination of channel features without expanding the spatio-temporal receptive field. In parallel, the LST component applies two consecutive $3\times 3\times 3$ convolutions with GELU activations, each followed by batch normalization, allowing the component to capture fine-grained spatio-temporal dependencies from local spatial and temporal neighborhoods. The outputs from CMix and LST are fused to combine channel-mixed cues with rich local spatio-temporal patterns. A second LST block is then applied to further refine temporal representations. Finally, the fused CMix + LST output is added residually to the output of the second LST block. The combined representation is subsequently passed through GELU activation, layer normalization, and an up-projection layer to restore the original feature dimension.
The TAPE output is then added residually to the output of the preceding Swin Transformer stage, completing one STeM block. The TAPE’s hierarchical design allows the network to learn rich spatio-temporal representations efficiently. Furthermore, the TAPE block is architecture-agnostic and can be integrated into other vision backbones for video-based tasks as will be explored further in Section IV-C.
Figure 3: MSTCN architecture with parallel 1D convolutional streams captures multi-scale temporal patterns, followed by 1D convolution, TLP layers, and a BiLSTM for long-term dependencies.
III-B Temporal Enhancement with MS-TCN and BiLSTM
Following the spatio-temporal encoding stage, we apply a temporal enhancement module to further refine the features with local and global temporal contexts. This module consists of an MS-TCN block followed by a BiLSTM. The MS-TCN is designed to capture temporal information at multiple scales while keeping the computational cost relatively low [alyami2025swin]. It employs several parallel 1D convolutional streams, each configured with a fixed kernel size, $3\times 1$, and different dilation rates to create diverse receptive fields. In total, the network contains $n+2$ branches, where $n$ corresponds to the number of dilations and the additional two branches are dedicated to reduce feature dimensionality. In our implementation, we set $n=4$, resulting in six parallel streams. The outputs of these parallel streams are concatenated and then processed through a sequence of 1D convolution and TLP layers. This structure enables the MS-TCN to effectively capture both fine-grained and coarse-grained temporal patterns, making it suitable for modeling local dependencies. The refined representation is then passed to a BiLSTM layer, which models bidirectional dependencies to capture long-term temporal context. Figure 3 illustrates the MS-TCN architecture.
III-C Training Objective
The training objective used in our framework uses a weighted sum of multiple loss functions designed to optimize both early and late stages of the model. We apply a CTC loss after the MS-TCN output ($\mathcal{L}_{\text{ConvCTC}}$) and after the final BiLSTM output ($\mathcal{L}_{\text{SeqCTC}}$). Both loss functions are computed against the ground-truth gloss sequence. To encourage consistency between early and late temporal representations, we include a distillation loss ($\mathcal{L}_{\text{Dist}}$)  [min2021visual] which aligns the intermediate MS-TCN logits with those produced by the BiLSTM. Additionally, two auxiliary regularization losses from the LiftPool operation within the MS-TCN are incorporated, namely the uncertainty loss ($\mathcal{L}_{\text{Cu}}$) and the peakiness loss ($\mathcal{L}_{\text{Cp}}$), which encourage temporal confidence and sparsity [hu2022temporal]. The final loss is computed as follows:
$\mathcal{L}=\lambda_{1}\mathcal{L}_{\text{ConvCTC}}+\lambda_{2}\mathcal{L}_{\text{SeqCTC}}+\lambda_{3}\mathcal{L}_{\text{Dist}}+\lambda_{4}\mathcal{L}_{\text{Cu}}+\lambda_{5}\mathcal{L}_{\text{Cp}},$
where $\lambda_{1}$ and $\lambda_{2}$ are set to 1.0 to equally supervise both CTC stages. We also assign a higher weight of $\lambda_{3}=25.0$ to the disillation loss to emphasize the importance of distillation, and use smaller values $\lambda_{4}=\lambda_{5}=0.001$ for the auxiliary LiftPool losses to act as regularizers without dominating the optimization.
IV Experiments
IV-A Experimental Setup
Datasets. Three benchmarked CSLR datasets have been used in this work to evaluate the proposed framework: PHOENIX14, PHOENIX14T, and CSL-Daily. PHOENIX14 is a German Sign Language (GSL) dataset compiled from televised weather reports. This dataset features a vocabulary of 1295 unique glosses across 6,841 annotated sentence-level samples, with an average of 11.5 glosses per sentence. The data is performed by nine signers and split into 5,672 training, 629 validation, and 540 test samples. PHOENIX14T extends PHOENIX14 with improved gloss annotations and corresponding spoken German translations. This dataset includes 8,247 videos with 1,085 glosses, and an average of 7.7 glosses per sentence. The dataset is split into 7,096 samples for training, 519 for validation, and 642 for testing. CSL-Daily is a large-scale Chinese Sign Language dataset comprising 6,588 unique sentences covering approximately 2,000 distinct glosses. Videos are recorded against a uniform white background, and each sentence contains around 7.2 glosses on average. The dataset is divided into 18,401 samples for training, 1,077 for validation, and 1,176 for testing.
Implementation Details. The proposed model was developed using PyTorch and trained on a dual-GPU setup consisting of two NVIDIA RTX A6000 cards, each with 48GB of VRAM. We experimented with three backbone configurations: Swintiny, Swinsmall and Swinbase. The tiny, small, and base variants differ primarily in model depth, embedding dimensions, and the number of attention heads across stages, with larger variants providing greater capacity at the cost of increased computation. All of these backbones were initialized from ImageNet-1K pretrained checkpoints.
Each frame in the input video was resized to $256\times 256$ pixels and random crops of size $224\times 224$ were extracted during training. At test time, frames were center-cropped to $224\times 224$. To increase temporal and spatial diversity, the training pipeline included horizontal flipping applied with a probability of 0.5 and temporal jittering within a 20% range. Training was carried out using an Adam optimizer, with a constant weight decay set to $1\times 10^{-4}$. For both PHOENIX14 and PHOENIX14T datasets, we trained models for 40 epochs, starting with a learning rate of $10^{-4}$ and decreasing it by a factor of 5 at four milestones: epochs 20, 25, 30, and 35. The CSL-Daily dataset was trained under a different schedule where the learning rate began at $5\times 10^{-5}$ and was halved at regular intervals, specifically, at epochs 25, 30, 35, 40, and 45 over a total of 50 epochs.
IV-B Comparison with Other Works
In this section, we present a comparative evaluation of our proposed model against existing approaches. Specifically, we compare our model with single-stream RGB-based methods, multi-stream frameworks, and multimodal systems. Following prior works [ahn2024slowfast, alyami2025swin, hu2024corrnet+, alyami2025swin], we employ the Word Error Rate (WER) as our primary evaluation metric. It is computed by calculating the number of insertions, deletions, and substitutions required to transform the predicted sequence into the ground-truth sequence. A lower WER indicates higher recognition accuracy.
RGB-Based Frameworks. Table I compares the performance of our USTM framework with prior single-stream RGB-based methods. As shown in the table, USTM variants demonstrate consistent improvements across all datasets. On Phoenix14, USTMtiny achieves WERs of 18.1 and 18.3 on the dev and test sets, respectively. USTMsmall achieves the same scores, while USTMbase improves to 17.4 on dev and 17.7 on test. Our smallest model remains highly competitive, trailing CorrNet+ by only 0.1 on both dev and test. USTMbase surpasses the previous SOTA CorrNet+ with WER of 18.0 on dev and 18.2 on test, setting a new SOTA on this dataset. Notably, USTMtiny, which uses the tiny Swin transformer backbone, outperforms Swin-MSTP that relies on the larger Swinsmall variant which underscores the efficacy of the proposed approach.
TABLE I: Comparison with single stream RGB based models.
Method Phoenix14 Phoenix14T CSL-Daily
Dev Test Dev Test Dev Test
FCN [cheng2020fully] 23.7 23.9 23.3 25.1
mLTSF-Net [xie2023multi] 22.9 23.0
VAC [min2021visual] 21.2 22.3
MSTNet [xie2023multi] 20.3 21.4
SEN [hu2023self] 19.5 21.0 19.3 20.7
SMKD [hao2021self] 20.8 21.0 20.8 22.4
TLP [hu2022temporal] 19.7 20.8 19.4 21.2
SSSLR [jang2023self] 20.9 20.7 20.5 22.3
CTCA [guo2023distilling] 19.5 20.3 19.3 20.3 31.3 29.4
ST-Transformer [cui2023spatial] 19.9 19.9
CorrNet [hu2023continuous] 18.8 19.4 18.9 20.5 30.6 30.1
CorrNet+ [hu2024corrnet+] 18.0 18.2 17.2 19.1 28.6 28.2
Swin-MSTP [alyami2025swin] 18.1 18.7 18.7 19.7 28.3 27.1
CLIP-SLA [alyami2025swin] 19.7 19.3 19.8 19.4 26.0 25.8
USTMtiny (Ours) 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) 18.1 18.3 17.4 19.3 27.7 26.4
USTMbase (Ours) 17.4 17.7 17.6 18.9 27.7 26.5
On Phoenix14T, USTMtiny obtained WERs of 17.8 and 19.4 on the dev and test sets, respectively, while USTMsmall reduces the WER to 17.4 on dev and 19.3 on test sets. The best performance is achieved by USTMbase with 17.6 on dev and 18.9 on test sets, marking the lowest test WER among all approaches. Compared to CorrNet+, our base variant closely matches the dev score and still improves the test score.
On CSL-Daily, USTMtiny obtains 27.8 on dev and 27.6 on test sets. USTMbase is slightly better with 27.7 on dev and 26.5 on test, while USTMsmall achieves the best performance with 27.7 on dev and 26.4 on test sets. Our USTM models significantly outperform CorrNet+ and Swin-MSTP. However, CLIP-SLA performs better on CSL-Daily likely due to its parameter-efficient fine-tuning strategy. By only updating LoRA modules within the pre-trained CLIP encoder, the model avoids overfitting to the training data, whereas USTM trains all parameters of the visual backbone, which may lead to slightly reduced generalization on this dataset.
Multi-Stream Frameworks. Table II compares the performance of our USTM framework with prior multi-stream frameworks that use multiple streams to capture varied feature representations. On Phoenix14, USTMbase surpasses all previous methods, including SlowFastSign, which achieved WERs of 18.0 and 18.3 on the dev and test sets, respectively. Our smaller variants achieve performance on par with SlowFastSign, demonstrating the efficiency of our single-stream design. On Phoenix14T, our models surpass previous methods on the dev split, while trailing the prior best on the test split by a WER of 0.2. On CSL-Daily, USTMsmall trails the previous best method by a margin of 2.2 on the dev split and 1.5 on the test split. The performance gap between our models and SlowFastSign can be attributed to two main factors. First, SlowFastSign is pretrained on the large-scale Kinetics-400 dataset, which provides a strong spatiotemporal prior, while our models are only initialized with ImageNet weights. Second, the SlowFast architecture is explicitly designed to model temporal dynamics at multiple scales throughout the network, which inherently benefits continuous sign language recognition. Despite these differences, our single-stream architecture remains highly competitive and demonstrates strong generalization across datasets without relying on multi-stream inputs or specialized temporal pretraining.
TABLE II: Comparison with multi-stream RGB-based models.
Method Phoenix14 Phoenix14T CSL-Daily
Dev Test Dev Test Dev Test
CMA [pu2020boosting] 23.9 24.0 24.1 24.3
AdaBrowse [hu2023adabrowse] 19.6 20.7 19.5 20.6 31.2 30.7
MSKA [ahn2024slowfast] 21.7 22.1 20.1 20.5 28.2 27.8
SlowFastSign [ahn2024slowfast] 18.0 18.3 17.7 18.7 25.5 24.9
USTMtiny (Ours) 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) 18.1 18.3 17.4 19.3 27.7 26.4
USTMbase (Ours) 17.4 17.7 17.6 18.9 27.7 26.5
Multi-Modal Frameworks. We also compare the performance of the USTM framework with other methods that utilize more than one modality. Table III compares the results of our framework with prior multi-modal methods. As shown in the table, most of the multi-modal methods use combine RGB with pose information. On the Phoenix14 dataset, USTMbase achieves the best results, surpassing all previous works. Notably, even our smallest variant, USTMtiny, outperforms TwoStream-SLR the SOTA multi-modal CSLR, achieving 18.1 in dev and 18.3 in test. USTMsmall also performs competitively with the same scores, matching or slightly improving over prior SOTA. For Phoenix14T, USTM improves over TwoStream-SLR with 17.7 in dev and 19.3 in test. On CSL-Daily, USTMsmall attains the lowest WER, surpassing TwoStream-SLR, with 33.1 in dev and 32.0 in test sets.
TABLE III: Comparison with multi-modal frameworks.
Method Modality Phoenix14 Phoenix14T CSL-Daily
Dev Test Dev Test Dev Test
DNF [cui2019deep] RGB+Optical Flow 23.1 22.8 32.8 32.4
SignBERT [zhou2021signbert] RGB+Pose 20.1 20.2 33.6 33.1
STMC [zhou2021spatial] RGB+Pose 20.6 21.5 19.6 21.0
C2SLR [zuo2022c2slr] RGB+Pose 20.5 20.4 20.2 20.4
MSTCSLRN [li2022multi] RGB+Gloss 22.8
CVT-SLR [zheng2023cvt] RGB+Pose 19.8 20.1 19.4 20.3
TwoStream-SLR [chen2022two] RGB+Pose 18.4 18.8 17.7 19.3 33.1 32.0
USTMtiny (Ours) RGB 18.1 18.3 17.8 19.4 27.8 27.6
USTMsmall (Ours) RGB 18.1 18.3 17.4 19.3 27.7 26.4
USTMbase (Ours) RGB 17.4 17.7 17.6 18.9 27.7 26.5
IV-C Ablation Studies
Figure 4: EigenCAM visualization of a sample from the Phoenix2014 dataset. The top row illustrates the EigenCAM of the USTM with a ViT backbone, while the bottom row illustrates the EigenCAM of the USTM with a Swin transformer backbone. The red color indicates the most attended regions.
To analyze the contribution of our framework’s components, we conducted a series of ablation studies on the Phoenix2014 dataset. These experiments evaluate the effect of the temporal adapter design, the spatio-temporal encoding strategy, and core architectural hyperparameters. In all experiments involving temporal adapters, we follow a consistent placement strategy unless stated otherwise. Specifically, for the Swin transformer backbone, the adapters are inserted after each of the four Swin transformer stages, whereas for the ViT backbone, they are integrated after every three transformer blocks.
Temporal Adapter Design. The TAPE adapter lies at the core of our USTM pipeline. We first evaluate whether incorporating adapters within the spatial backbone is beneficial to the overall performance of the proposed framework. We adopt STAdapter [pan2022st] as a baseline spatiotemporal adapter and compare it with the proposed TAPE adapter. In the temporal enhancement stage, we employed a simple Conv1D + BiLSTM module instead of MS-TCN + BiLSTM, ensuring that the effect of MS-TCN does not confound the comparison. As shown in Table IV, interleaving STAdapter blocks within the spatial backbone substantially improves performance. The effect is most apparent with the ViT backbone, where WER drops from 43.2 to 22.3 on dev and from 43.3 to 22.1 on test. A similar trend is observed with the Swin transformer backbone, where WER decreases by 0.6 on dev and 0.8 on test. Replacing STAdapter with TAPE yields further improvements, reducing WER on ViT to 20.2 and 20.7, and on Swinsmall to 18.4 and 19.1 on dev and test, respectively.
TABLE IV: Effect of spatiotemporal adapters.
Backbone Adapter Dev Test
ViT - 43.2 43.3
ViT STAdapter 22.3 22.1
ViT TAPE 20.2 20.7
Swinsmall - 19.9 19.9
Swinsmall STAdapter 19.3 19.1
Swinsmall TAPE 18.4 19.1
Temporal Adapter Insertion Frequency. For deeper architectures, such as ViT-B16, which consists of 12 layers compared to only 4 stages in Swin, we studied how frequently TAPE adapters should be inserted. To control this, we introduce a hyperparameter $\alpha$ that defines the interval between two TAPE adapters, where smaller $\alpha$ values correspond to denser insertion. As shown in Table V, inserting TAPE after every single layer ($\alpha=1$) or too sparsely ($\alpha=3$) results in weaker performance. The best performance is achieved with $\alpha=2$, suggesting that moderate temporal injection offers the right balance between model capacity and temporal granularity, allowing effective exploitation of local temporal structure without oversaturating the backbone.
TABLE V: Effect of varying adapter insertion interval $\alpha$ in ViT-B16.
Adapter Interval $\alpha$ Dev Test
1 22.8 22.6
2 20.4 20.0
3 20.8 20.7
USTM Components. We first analyze the choice of backbone to study its impact on the USTM pipeline. ViT-B16 and Swinsmall were compared under minimal configurations using only a simple Conv1D and BiLSTM in the temporal enhancement stage. As shown in Table VI, ViT-B16 performs poorly with a WER of 43.2 on dev and 43.3 on test, while Swinsmall achieves a WER of 19.8 on dev and 19.9 on test. The superiority of Swin over ViT is further evident in the EigenCam [muhammad2020eigen] visualizations in Figure 4. The Swin-based model focuses clearly on key regions, such as the hands and mouth, producing smooth and semantically meaningful attention maps. In contrast, the ViT-based model attends to irrelevant areas like the torso, with noisy and fragmented attention. This highlights Swin’s strength in capturing spatially dense cues through its hierarchical and windowed attention, reinforcing its effectiveness for CSLR tasks.
TABLE VI: Ablation study of USTM core components, evaluating the effects of backbone selection, TAPE, and MS-TCN.
Backbone TAPE MS-TCN Dev Test
ViT-B16 Swinsmall
43.2 43.3
19.3 19.6
20.2 20.7
20.8 20.7
19.8 19.9
18.1 18.7
18.4 19.1
17.9 17.6
Next, we investigate the interplay between the TAPE adapter and MS-TCN as temporal modeling components. As previously established, TAPE alone improves performance on both ViT-B16 and Swin transformer backbones. Incorporating MS-TCN further refines temporal modeling. For ViT-B16, MS-TCN alone reduces WER to 19.3 and 19.6 on dev and test respectively. However, MS-TCN combined with TAPE increase the WER, indicating some redundancy in this configuration. For Swinsmall, however, combining MS-TCN with TAPE yields the best overall performance of 17.9 on dev and 17.6 on test. These findings indicate that MS-TCN and TAPE interact differently across backbones, showing complementary behavior for Swin but partial redundancy for ViT-B16. This suggests that their combined benefit is architecture-dependent.
V Conclusion
In this work, we introduced USTM, a unified spatio-temporal modeling framework for CSLR. By integrating hierarchical visual backbones, such as the Swin Transformer, with the TAPE adapter and MS-TCN temporal modules, USTM effectively captures fine-grained spatial details and multi-scale temporal dependencies in sign language videos. Our design allows temporal information to be directly injected into the visual backbone while simultaneously modeling long-range motion patterns, resulting in robust and semantically meaningful representations. Extensive experiments on PHOENIX14, PHOENIX14T, and CSL-Daily datasets demonstrate that USTM consistently outperforms prior single-stream, multi-stream, and multi-modal CSLR approaches, establishing new SOTA performance. Ablation studies further confirm the importance of adapter placement, temporal modeling, and backbone selection, while EigenCAM quantitative analysis illustrated the model’s capability in capturing critical regions on the signer, such as the hands and face. Despite its strong performance, USTM relies on a relatively large Swin transformer backbone, which can pose computational challenges for real-time deployment. Future work may explore model compression, efficient temporal adapters, and multi-modal extensions to enhance generalization and enable deployment in resource-constrained environments.
Acknowledgments
The authors would like to acknowledge the support provided by King Fahd University of Petroleum & Minerals (KFUPM) for funding this work through project number ISP24226.